OpenStudy (anonymous):

What is the nth term of the geometric sequence 4, 8, 16, 32, ... ?

OpenStudy (anonymous):

`a_(n)=2(4)^(n-1)` `a_(n)=4(2)^(n-1)` `a_(n)=2(n)^4` `a_n=4(n)^2`

OpenStudy (science0229):

The first term is 4, and it's multiplying by 2. So for the second term, it would be 4*2 For the third term, it would be 4*2^2. For the nth term, it would be \[4*2^{n-1}\]

OpenStudy (anonymous):

so what would be my answer from my option that i wrote